OpenStudy (anonymous):

You roll a fair six sided die twice. What is the probability that the die will land with the same side facing up both times? 1. 1⁄6 2. 1⁄12 3. 2⁄36 4. 1⁄36 5. 0

OpenStudy (lgbasallote):

probability that it will land on one certain number is \[\frac{1}{6}\]probability it will roll to that number again is \[\frac{1}{6} \times \frac{1}{6}\] got that?

OpenStudy (lgbasallote):

are you there @anum93 ?

OpenStudy (lgbasallote):

do you get the solution and how to arrive with the answer??

OpenStudy (anonymous):

you sure becoz some where i see the answer 1/6

OpenStudy (anonymous):

you sure becoz some where i see the answer 1/6

OpenStudy (lgbasallote):

considering as i got some medals for my answers...i assume they agree with me so im sure ;)